OpenStudy (anonymous):

Gaseous diffusion is used to enrich natural uranium in U-235. Use Graham’s law (see feedback) to calculate the ratio of enrichment of U-235 by a single diffusion of UF6. The atomic masses of the two principal isotopes of uranium are 235.0493 and 238.0508 for U-235 and U-238, respectively, and that of F is 18.9984. Use 4 decimal places.

OpenStudy (aaronq):

Graham's law is \(\large \sf \dfrac{rate_1}{rate_2}=\sqrt{\dfrac{M_2}{M_1}}\) we need the molar masses, \(M_1=M_{U-235}= 235.0493+6* 18.9984=349.0397~g/mol\) \(M_2=M_{U-238}= 238.0508 +6* 18.9984=352.0412~g/mol\) \(\large \sf \dfrac{rate_{U-235}}{rate_{U-238}}=\sqrt{\dfrac{352.0412}{349.0397}}\) \(\large \sf \dfrac{rate_{U-235}}{rate_{U-238}}=1.0042904501840236059=1.0043\)